The product and sum of three consecutive terms in a Geometric Progression are 3375 and 65 respectively. Find the difference between the largest and smallest terms of the three numbers
Answer by stanbon(75887)

The product and sum of three consecutive terms in a Geometric Progression are 3375 and 65 respectively. Find the difference between the largest and smallest terms of the three numbers
-----
Product: (a*ar*ar^2) = 3375
(ar)^3 = 3375
ar = 15
-----
Sum::::: (a + ar + ar^2) = 65
(a + ar + ar*r) = 65
----
Substitute for ar in the sum to get
(a + 15 + 15r) = 65
------
Two Equations::
ar = 15
a + 15r = 50
-----
Substitute for "a" and solve for "r":
(50-15r)r = 15
----
15r^2 - 50r + 15 = 0
----
3r^2 - 10r + 3 = 0
Factor:
3r^2-9r-r+3 = 0
3r(r-3)-(r-3) = 0
(r-3)(3x-1) = 0
===
r = 3 or r = 1/3
----
If r = 3, a = 5 : Then ar^2 = 45:: Then difference = 45-5 = 40
If r = 1/3, a = 45: Then ar^2 = 5:: Then diffeerence = 5-45 = -40
=====================
Cheers,
Stan H.
=====================
-----


-----